# Distinguishing radiation mechanisms and particle populations in blazar jets through long-term multi-band monitoring with RINGO3 and Fermi

### v1

@@ Appendix A False correlations in colour-magnitude analysis
When correlation analysis is performed on datasets where the same data are used on two axes, they contain a common uncertainty which may produce a false correlation due to correlated errors. To demonstrate this we use the example of a simple colour magnitude correlation. $b$ vs $b-r$ and $r$ vs $b-r$ plots were made with RINGO3 data to compare with the analysis made by Gupta et al. (2016). The correlations suggested that sources were both bluer- and redder-when-brighter, depending on the chosen x-axis (b* magnitude and r* magnitude respectively). This highlights the necessity to use two independent data sets when performing correlation statistics (i.e. both data sets must have independent uncertainties)
To explore this property independently from RINGO3 data, random magnitude values (b and r) were generated using Monte Carlo methods, along with randomly generated error values, and plotted against each other in the form b vs b-r and r vs b-r. A linear regression was fitted to each set of data points and this showed the preference for x-axis-dependent correlations. Fig. 5 shows an example of 3 repetitions of this process with one hundred values centred on 15, and error values centred on 2. These trends highlight the necessity to have truly independent data for correlation analysis. For this reason the optical spectral index (a similar measurement to colour) calculated in this work uses the r* and b* flux only, so as to not include a common uncertainty on both the x and y axes when correlated against the g*-band flux.
@@ Appendix B Figures and Tables
Source Type r* c r* p g* c g* p b* c b* p Points
IC 310 HSP - - - - - - 206
1ES 1011+496 HSP -0.113 0.144 -0.088 0.252 -0.107 0.165 170
MRK 421 HSP - - - - - - 281
MRK 180 HSP 0.117 0.126 0.110 0.152 0.137 0.074 171
PG 1218+304 HSP 0.043 0.653 0.002 0.984 0.027 0.777 110
1ES 1426+428 HSP - - - - - - 219
PG 1553+113 HSP 0.482 2.04×$\times 10^{-16}$ 0.480 3.00×$\times 10^{-16}$ 0.471 1.11×$\times 10^{-15}$ 258
MRK 501 HSP 0.366 1.21×$\times 10^{-11}$ 0.391 3.14×$\times 10^{-13}$ 0.365 1.42×$\times 10^{-11}$ 322
1ES 1959+650 HSP -0.186 0.196 -0.145 0.315 -0.140 0.332 50
3C 66A ISP 0.438 6.02×$\times 10^{-15}$ 0.419 1.21×$\times 10^{-13}$ 0.439 5.22×$\times 10^{-15}$ 288
S5 0716+714 ISP 0.705 2.99×$\times 10^{-44}$ 0.710 3.92×$\times 10^{-45}$ 0.720 6.29×$\times 10^{-47}$ 286
ON 231 ISP -0.129 0.247 -0.179 0.108 -0.128 0.251 82
A0 0235+164 LSP 0.674 2.94×$\times 10^{-10}$ 0.674 3.08×$\times 10^{-10}$ 0.660 9.20×$\times 10^{-10}$ 68
TXS 0506+056 LSP 0.569 0.034 0.670 8.78×$\times 10^{-3}$ 0.670 8.78×$\times 10^{-3}$ 14
OJ 287 LSP 0.266 4.43×$\times 10^{-10}$ 0.254 2.62×$\times 10^{-9}$ 0.229 8.76×$\times 10^{-8}$ 534
S4 0954+65 LSP 0.472 2.60×$\times 10^{-12}$ 0.476 1.50×$\times 10^{-12}$ 0.465 6.06×$\times 10^{-12}$ 197
4C 09.57 LSP 0.458 9.20×$\times 10^{-8}$ 0.452 1.41×$\times 10^{-7}$ 0.454 1.20×$\times 10^{-7}$ 124
BL Lac LSP 0.661 8.46×$\times 10^{-50}$ 0.674 2.44×$\times 10^{-52}$ 0.681 8.68×$\times 10^{-54}$ 385
PKS 0502+049 LSP (FSRQ) 0.824 7.75×$\times 10^{-30}$ 0.824 6.12×$\times 10^{-30}$ 0.829 1.36×$\times 10^{-30}$ 116
PKS 0736+01 LSP (FSRQ) 0.844 2.05×$\times 10^{-35}$ 0.837 3.37×$\times 10^{-34}$ 0.852 1.31×$\times 10^{-36}$ 126
PKS 1222+216 LSP (FSRQ) 0.438 4.74×$\times 10^{-5}$ 0.405 1.98×$\times 10^{-4}$ 0.398 2.60×$\times 10^{-4}$ 80
3C 279 LSP (FSRQ) 0.465 $\times 10^{-9}$ 0.478 2.94×$\times 10^{-9}$ 0.464 9.74×$\times 10^{-9}$ 138
PKS 1510-089 LSP (FSRQ) 0.564 2.49×$\times 10^{-21}$ 0.559 7.33×$\times 10^{-21}$ 0.521 6.35×$\times 10^{-18}$ 237
OS 319 LSP (FSRQ) 0.217 0.049 0.353 1.06×$\times 10^{-3}$ 0.264 0.016 83
PKS B1622-297 LSP (FSRQ) 0.272 0.012 0.222 0.042 0.231 0.035 84
4C +38.41 LSP (FSRQ) 0.668 6.13×$\times 10^{-13}$ 0.639 1.21×$\times 10^{-11}$ 0.644 7.50×$\times 10^{-12}$ 90
3C 345 LSP (FSRQ) 0.089 0.381 0.035 0.729 0.024 0.815 99
PKS B1730-130 LSP (FSRQ) 0.026 0.805 0.057 0.593 0.071 0.506 90
3C 446 LSP (FSRQ) 0.318 4.52×$\times 10^{-3}$ 0.332 2.99×$\times 10^{-3}$ 0.261 0.021 78
4C 11.69 LSP (FSRQ) 0.761 2.36×$\times 10^{-42}$ 0.761 3.24×$\times 10^{-42}$ 0.760 4.01×$\times 10^{-42}$ 217
3C 454.3 LSP (FSRQ) 0.891 4.72×$\times 10^{-37}$ 0.887 2.50×$\times 10^{-36}$ 0.869 3.63×$\times 10^{-33}$ 105
@@
Table 5: Spearman correlation coefficients for optical flux vs Fermi $\gamma$-ray flux, where $c$ is the correlation coefficient and $p$ is the corresponding p value. The number of optical data points (per camera) used in these correlation calculations is also shown. Note there are no Fermi data available for IC 310 or 1ES 1426+428.
Source Type Fit c p Trend $\alpha_{\mathrm{av}}$ Points
IC 310 HSP linear -0.288 2.72×$\times 10^{-5}$ BWB 1.134 206
1ES 1011+496 HSP linear -0.251 9.51×$\times 10^{-4}$ BWB -0.233 170
MRK 421 HSP - - - - - 281
MRK 180 HSP linear -0.527 1.29×$\times 10^{-13}$ BWB 0.578 171
PG 1218+304 HSP log -0.735 6.11×$\times 10^{-20}$ BSWB 0.091 110
1ES 1426+428 HSP linear -0.266 6.84×$\times 10^{-5}$ BWB 0.426 219
PG 1553+113 HSP linear -0.378 3.51×$\times 10^{-10}$ BWB -0.211 258
MRK 501 HSP linear -0.527 2.32×$\times 10^{-24}$ BWB 0.416 322
1ES 1959+650 HSP linear -0.088 0.543 - 0.093 50
3C 66A ISP linear -0.195 9.01×$\times 10^{-4}$ BWB 0.065 288
S5 0716+714 ISP log -0.639 3.38×$\times 10^{-34}$ BSWB 0.123 286
ON 231 ISP log -0.487 3.49×$\times 10^{-6}$ BSWB 0.336 82
A0 0235+164 LSP linear 0.164 0.182 - 1.431 68
TXS 0506+056 LSP linear -0.881 3.11×$\times 10^{-5}$ BWB 0.271 14
OJ 287 LSP log -0.494 2.92×$\times 10^{-34}$ BSWB 0.149 534
S4 0954+65 LSP log -0.598 1.74×$\times 10^{-20}$ BSWB 0.576 197
4C 09.57 LSP linear -0.421 1.11×$\times 10^{-6}$ BWB 0.750 124
BL Lac LSP log -0.637 3.51×$\times 10^{-45}$ BSWB 0.931 385
PKS 0502+049 LSP (FSRQ) log 0.825 5.64×$\times 10^{-30}$ RSWB -0.172 116
PKS 0736+01 LSP (FSRQ) log -0.816 2.35×$\times 10^{-31}$ BSWB 0.774 126
PKS 1222+216 LSP (FSRQ) log 0.769 7.98×$\times 10^{-17}$ RSWB -0.154 80
3C 279 LSP (FSRQ) log -0.374 6.18×$\times 10^{-6}$ BSWB 0.524 138
PKS 1510-089 LSP (FSRQ) log 0.566 1.76×$\times 10^{-21}$ RSWB 0.136 237
OS 319 LSP (FSRQ) linear 0.227 0.039 RWB -0.290 83
PKS B1622-297 LSP (FSRQ) log 0.476 4.85×$\times 10^{-6}$ RSWB 0.599 84
4C +38.41 LSP (FSRQ) log 0.261 0.013 RSWB 0.048 90
3C 345 LSP (FSRQ) linear -0.210 0.037 BWB 0.243 99
PKS B1730-130 LSP (FSRQ) log 0.426 2.78×$\times 10^{-5}$ RSWB 1.077 90
3C 446 LSP (FSRQ) linear 0.243 0.032 RWB 0.120 78
4C 11.69 LSP (FSRQ) log 0.723 2.16×$\times 10^{-36}$ RSWB 0.469 217
3C 454.3 LSP (FSRQ) log 0.953 2.87×$\times 10^{-55}$ RSWB -0.167 105
@@
Table 6: Correlation strengths for the spectral index vs g*-band flux correlations. The source name and subclass are given in columns one and two. The best fit determined as given by AIC and BIC values is shown in column three. Columns four and five give the Spearman rank strength and significance correlation coefficients after having linearised the dataset if better fitted with a logarithmic curve. Column six gives the colour trend of the object given the fit and Spearman rank coefficients. Column seven gives the average spectral index, and column 8 gives the number of data points.
@@
Table 7: As Table 6 but for spectral index vs $\gamma$-ray flux.
@@
Table 8: DCF peak lag values with correlation strengths after computing on the optical and $\gamma$-ray fluxes. A negative lag implies the $\gamma$-ray emission is leading the optical.

### v2

@@ Appendix A False correlations in colour-magnitude analysis
When correlation analysis is performed on datasets where the same data are used on two axes, they contain a common uncertainty which may produce a false correlation due to correlated errors. To demonstrate this we use the example of a simple colour magnitude correlation. $b$ vs. $b-r$ and $r$ vs. $b-r$ plots were made with RINGO3 data to compare with the analysis made by Gupta et al. (2016). The correlations suggested that sources were both bluer- and redder-when-brighter, depending on the chosen x-axis (b* magnitude and r* magnitude respectively). This highlights the necessity to use two independent data sets when performing correlation statistics (i.e. both data sets must have independent uncertainties)
To explore this property independently from RINGO3 data, random magnitude values ($b$ and $r$) were generated using Monte Carlo methods, along with randomly generated error values, and plotted against each other in the form $b$ vs. $b$-$r$ and $r$ vs. $b$-$r$. A linear regression was fitted to each set of data points and this showed the preference for x-axis-dependent correlations. Fig. 5 shows an example of 3 repetitions of this process with one hundred values centred on 15, and error values centred on 2. These trends highlight the necessity to have truly independent data for correlation analysis. For this reason the optical spectral index (a similar measurement to colour) calculated in this work uses the r* and b* flux only, so as to not include a common uncertainty on both the x and y axes when correlated against the g*-band flux.
@@ Appendix B Figures and Tables
Table 5: Spearman correlation coefficients for optical flux vs. Fermi $\gamma$-ray flux, where $c$ is the correlation coefficient and $p$ is the corresponding significance coefficient
. The number of optical data points (per camera) used in these correlation calculations is also shown. Note there are no Fermi data available for IC 310 or 1ES 1426+428. Source Type r* c r* p g* c g* p b* c b* p Points IC 310 HSP - - - - - - 206 1ES 1011+496 HSP -0.113 0.144 -0.088 0.252 -0.107 0.165 170 MRK 421 HSP - - - - - - 281 MRK 180 HSP 0.117 0.126 0.110 0.152 0.137 0.074 171 PG 1218+304 HSP 0.043 0.653 0.002 0.984 0.027 0.777 110 1ES 1426+428 HSP - - - - - - 219 PG 1553+113 HSP 0.482 2.04×$\times 10^{-16}$ 0.480 3.00×$\times 10^{-16}$ 0.471 1.11×$\times 10^{-15}$ 258 MRK 501 HSP 0.366 1.21×$\times 10^{-11}$ 0.391 3.14×$\times 10^{-13}$ 0.365 1.42×$\times 10^{-11}$ 322 1ES 1959+650 HSP -0.186 0.196 -0.145 0.315 -0.140 0.332 50 3C 66A ISP 0.438 6.02×$\times 10^{-15}$ 0.419 1.21×$\times 10^{-13}$ 0.439 5.22×$\times 10^{-15}$ 288 S5 0716+714 ISP 0.705 2.99×$\times 10^{-44}$ 0.710 3.92×$\times 10^{-45}$ 0.720 6.29×$\times 10^{-47}$ 286 ON 231 ISP -0.129 0.247 -0.179 0.108 -0.128 0.251 82 A0 0235+164 LSP 0.674 2.94×$\times 10^{-10}$ 0.674 3.08×$\times 10^{-10}$ 0.660 9.20×$\times 10^{-10}$ 68 TXS 0506+056 LSP 0.569 0.034 0.670 8.78×$\times 10^{-3}$ 0.670 8.78×$\times 10^{-3}$ 14 OJ 287 LSP 0.266 4.43×$\times 10^{-10}$ 0.254 2.62×$\times 10^{-9}$ 0.229 8.76×$\times 10^{-8}$ 534 S4 0954+65 LSP 0.472 2.60×$\times 10^{-12}$ 0.476 1.50×$\times 10^{-12}$ 0.465 6.06×$\times 10^{-12}$ 197 4C 09.57 LSP 0.458 9.20×$\times 10^{-8}$ 0.452 1.41×$\times 10^{-7}$ 0.454 1.20×$\times 10^{-7}$ 124 BL Lac LSP 0.661 8.46×$\times 10^{-50}$ 0.674 2.44×$\times 10^{-52}$ 0.681 8.68×$\times 10^{-54}$ 385 PKS 0502+049 LSP (FSRQ) 0.824 7.75×$\times 10^{-30}$ 0.824 6.12×$\times 10^{-30}$ 0.829 1.36×$\times 10^{-30}$ 116 PKS 0736+01 LSP (FSRQ) 0.844 2.05×$\times 10^{-35}$ 0.837 3.37×$\times 10^{-34}$ 0.852 1.31×$\times 10^{-36}$ 126 PKS 1222+216 LSP (FSRQ) 0.438 4.74×$\times 10^{-5}$ 0.405 1.98×$\times 10^{-4}$ 0.398 2.60×$\times 10^{-4}$ 80 3C 279 LSP (FSRQ) 0.465 $\times 10^{-9}$ 0.478 2.94×$\times 10^{-9}$ 0.464 9.74×$\times 10^{-9}$ 138 PKS 1510-089 LSP (FSRQ) 0.564 2.49×$\times 10^{-21}$ 0.559 7.33×$\times 10^{-21}$ 0.521 6.35×$\times 10^{-18}$ 237 OS 319 LSP (FSRQ) 0.217 0.049 0.353 1.06×$\times 10^{-3}$ 0.264 0.016 83 PKS B1622-297 LSP (FSRQ) 0.272 0.012 0.222 0.042 0.231 0.035 84 4C +38.41 LSP (FSRQ) 0.668 6.13×$\times 10^{-13}$ 0.639 1.21×$\times 10^{-11}$ 0.644 7.50×$\times 10^{-12}$ 90 3C 345 LSP (FSRQ) 0.089 0.381 0.035 0.729 0.024 0.815 99 PKS B1730-130 LSP (FSRQ) 0.026 0.805 0.057 0.593 0.071 0.506 90 3C 446 LSP (FSRQ) 0.318 4.52×$\times 10^{-3}$ 0.332 2.99×$\times 10^{-3}$ 0.261 0.021 78 4C 11.69 LSP (FSRQ) 0.761 2.36×$\times 10^{-42}$ 0.761 3.24×$\times 10^{-42}$ 0.760 4.01×$\times 10^{-42}$ 217 3C 454.3 LSP (FSRQ) 0.891 4.72×$\times 10^{-37}$ 0.887 2.50×$\times 10^{-36}$ 0.869 3.63×$\times 10^{-33}$ 105
Table 6: Correlation strengths for the spectral index vs. g*-band flux correlations. The source name and subclass are given in columns one and two. The best fit functional form
@@
determined by AIC and BIC values is shown in column three. Columns four and five give the Spearman rank strength and significance correlation coefficients after having linearised the dataset if better fitted with a logarithmic curve. Column six gives the colour trend of the object given the fit and Spearman rank coefficients. Column seven gives the average spectral index, and column 8 gives the number of data points. Source Type Fit $c$ $p$ Trend $\alpha_{\mathrm{av}}$ Points IC 310 HSP linear -0.288 2.72×$\times 10^{-5}$ BWB 1.134 206 1ES 1011+496 HSP linear -0.251 9.51×$\times 10^{-4}$ BWB -0.233 170 MRK 421 HSP - - - - - 281 MRK 180 HSP linear -0.527 1.29×$\times 10^{-13}$ BWB 0.578 171 PG 1218+304 HSP log -0.735 6.11×$\times 10^{-20}$ BSWB 0.091 110 1ES 1426+428 HSP linear -0.266 6.84×$\times 10^{-5}$ BWB 0.426 219 PG 1553+113 HSP linear -0.378 3.51×$\times 10^{-10}$ BWB -0.211 258 MRK 501 HSP linear -0.527 2.32×$\times 10^{-24}$ BWB 0.416 322 1ES 1959+650 HSP linear -0.088 0.543 - 0.093 50 3C 66A ISP linear -0.195 9.01×$\times 10^{-4}$ BWB 0.065 288 S5 0716+714 ISP log -0.639 3.38×$\times 10^{-34}$ BSWB 0.123 286 ON 231 ISP log -0.487 3.49×$\times 10^{-6}$ BSWB 0.336 82 A0 0235+164 LSP linear 0.164 0.182 - 1.431 68 TXS 0506+056 LSP linear -0.881 3.11×$\times 10^{-5}$ BWB 0.271 14 OJ 287 LSP log -0.494 2.92×$\times 10^{-34}$ BSWB 0.149 534 S4 0954+65 LSP log -0.598 1.74×$\times 10^{-20}$ BSWB 0.576 197 4C 09.57 LSP linear -0.421 1.11×$\times 10^{-6}$ BWB 0.750 124 BL Lac LSP log -0.637 3.51×$\times 10^{-45}$ BSWB 0.931 385 PKS 0502+049 LSP (FSRQ) log 0.825 5.64×$\times 10^{-30}$ RSWB -0.172 116 PKS 0736+01 LSP (FSRQ) log -0.816 2.35×$\times 10^{-31}$ BSWB 0.774 126 PKS 1222+216 LSP (FSRQ) log 0.769 7.98×$\times 10^{-17}$ RSWB -0.154 80 3C 279 LSP (FSRQ) log -0.374 6.18×$\times 10^{-6}$ BSWB 0.524 138 PKS 1510-089 LSP (FSRQ) log 0.566 1.76×$\times 10^{-21}$ RSWB 0.136 237 OS 319 LSP (FSRQ) linear 0.227 0.039 RWB -0.290 83 PKS B1622-297 LSP (FSRQ) log 0.476 4.85×$\times 10^{-6}$ RSWB 0.599 84 4C +38.41 LSP (FSRQ) log 0.261 0.013 RSWB 0.048 90 3C 345 LSP (FSRQ) linear -0.210 0.037 BWB 0.243 99 PKS B1730-130 LSP (FSRQ) log 0.426 2.78×$\times 10^{-5}$ RSWB 1.077 90 3C 446 LSP (FSRQ) linear 0.243 0.032 RWB 0.120 78 4C 11.69 LSP (FSRQ) log 0.723 2.16×$\times 10^{-36}$ RSWB 0.469 217 3C 454.3 LSP (FSRQ) log 0.953 2.87×$\times 10^{-55}$ RSWB -0.167 105
Table 7: As Table 6 but for spectral index vs. $\gamma$-ray flux.
@@
Table 8: DCF peak lag values with correlation strengths after computing on the optical and $\gamma$-ray fluxes. A negative lag implies the $\gamma$-ray emission is leading the optical. An asterisk in the $3\sigma$ lag column suggests the potential significance of the detected lag, warranting further analysis.